OpenStudy (jarp0120):

hi guys can you help me please? Water flows out of a hemispherical tank at a rate which is 4 times the square root of its depth. If the radius of the tank is 8 ft, how fast is the water level falling when the water is 4 ft deep? Hint: Use the formula for the volume of a spherical segment, i.e. V=1/3 π h^2 (3r-h). wher the answer is 16 pi

OpenStudy (jarp0120):

\[h= 4 ft, r=8ft \frac{ dh }{ dt }= ?\]

zepdrix (zepdrix):

Hmm I think there is another piece of information you need before you do any differentiation :) `Water flows out of a hemispherical tank at a rate` v' which is `4 times` the `square root of its depth`. 4*sqrt(h) So we have: \(\large\rm v'=-4\sqrt{h}\) ya? :o

OpenStudy (baru):

@zepdrix continue please ...i seem to be going wrong somewhere

OpenStudy (baru):

well... i tried this: dV= pi*r^2 dh (where r is the radius that corresponds to height h)

OpenStudy (baru):

substitute for dV in V' \[\pi r^2 \frac{ dh }{ dt }= -4\sqrt{h}\]

zepdrix (zepdrix):

They gave us a nice formula to use, \(\large\rm v=\frac{\pi}{3}h^2(3r-h)\) This is the volume of the filled portion of the hemisphere. I guess we want to differentiate this formula with respect to time. Before we can do that though, we would like to write the r in terms of h. So I think we need to use another formula, let's call it v tilde, \(\large\rm \tilde{v}=\frac{2\pi}{3}r^3\) Oh no no no, scratch that... it's a lot simpler than that. r is held constant, it is r=8. So we don't even need to worry about any r' showing up. AHHH my bad.

zepdrix (zepdrix):

I dunno what you did baru :c sorry kinda confused by your formula thing

OpenStudy (baru):

do you get 16pi by ur method?

zepdrix (zepdrix):

\[\large\rm v=\frac{\pi}{3}h^2(3r-h)\]\[\large\rm v=\frac{\pi}{3}h^2(24-h)\]And then differentiate with respect to time, ya?\[\large\rm v'=\frac{\pi}{3}\left(h^2\right)'(24-h)+\frac{\pi}{3}h^2(24-h)'\]Product rule? :o

zepdrix (zepdrix):

Oh she provided the answer, nice. Hope so! Let's find out :d

zepdrix (zepdrix):

\[\large\rm v'=\frac{2\pi}{3}2h(24-h)h'-\frac{\pi}{3}h^2h'\]\[\large\rm v'=\left(\frac{2\pi}{3}2h(24-h)-\frac{\pi}{3}h^2\right)h'\]

zepdrix (zepdrix):

At \(\large\rm h=4\) and \(\large\rm v'=-4\sqrt{h}=-8\) we have,\[\large\rm -8=\left(\frac{2\pi}{3}2\cdot4(24-4)-\frac{\pi}{3}16\right)h'\]\[\large\rm -\frac{1}{2}=\left(\frac{\pi}{3}(20)-\frac{\pi}{3}\right)h'\]\[\large\rm -\frac{1}{2}=\left(\frac{19\pi}{3}\right)h'\]Hmm darn, no that's not leading to the answer which was provided :(\[\large\rm h'=-\frac{3}{38\pi}\]

zepdrix (zepdrix):

Hmmm

OpenStudy (baru):

i get -1/6pi .... in either case... it looks like 'pi' has to be in the denominator

zepdrix (zepdrix):

Well I guess she ran off to bed or something XD I should do the same heh

OpenStudy (baru):

XD

OpenStudy (jarp0120):

hi @baru and @zepdrix i got -1/6pi is it okay if the answer given by the book is wrong?

OpenStudy (jarp0120):

oh and thanks for your help

OpenStudy (baru):

it's possible that the text book has got it wrong...check with your teachers/classmates and let us know :)